ii. A mass of gas is stored in a cylinder. The gas is stored at a pressure of 1.2 MPa and the cylinder has a volume of 0.03 m3 and temperature of 100 oC , Find the mass of the gas.

The characteristic gas constant for air is 300 Jkg−1K−1.


1
Expert's answer
2021-05-05T07:41:13-0400

Given:

Pressure (P)=1.2 MPa = 1.2 x106 Pa

Volume (V)= 0.03 m3

Temperature (T)= 100 oC=373 K

The characteristic gas constant for air (R) = 300 Jkg−1K−1.


To Find :

Mass of the gas(M)


Solution :

Assuming air as an ideal gas , It obeys ideal gas equation P.V=M RT

Mass (M) of air is given by

"M=\frac{P\times V}{R\times T}"

"M=\frac{1.2\times 10^6\times 0.03}{300\times 373}=" 0.322 kg